An object moves along the x-axis directly from x1 = –12 cm to x2 = +12 cm, where it changes direction abruptly and moves directly to x3 = +6 cm, all in 10 s.
what is the average speed of the object?
1
Expert's answer
2021-09-29T09:53:57-0400

The average speed is the total distance covered over total time:


"s=\frac DT=\frac{12+12+6}{10}=3\text{ m/s}."